OpenStudy (laddiusmaximus):

@zepdrix The displacement in meters of an object moving in a straight line is given by s=1+2t+(1/4)t^2 where t is measured in seconds. (i) [1,3] (ii) [1,2] (iii) [1,1.5] (iv) [1,1.1] b) find the instantaneous velocity when t=1

OpenStudy (laddiusmaximus):

i use the formula lim h->0 f(a+h)-f(a)/h right?

zepdrix (zepdrix):

Ummm without the limit, yes. The limit tells us the "instantaneous rate of change", the same as the derivative. In this problem they want the "average rate of change", so we'll just use the difference quotient (the thingy in your limit problem without the limit attached to it).

OpenStudy (laddiusmaximus):

ok then that is 1/2a+1/4h

zepdrix (zepdrix):

Umm it's a little confusing the way they wrote it all abbreviated like that... I assume (i) is asking for the "average velocity from 1 to 3 seconds".

OpenStudy (laddiusmaximus):

I guess. it confused me

zepdrix (zepdrix):

\[\large v_{ave}=\frac{ s(3)-s(1) }{ 3-1 }\] This is how we'll find the average velocity between 3 and 1. I'm not sure what I was rambling on about with that whole limit thing XD lol

zepdrix (zepdrix):

Maybe you remember this idea from algebra. \[\large m=\frac{ y_2-y_1 }{ x_2-x_1 }\] We're doing the same thing with our problem, but we're using the fancy function notation and letting x and y be time and displacement.

OpenStudy (laddiusmaximus):

oh so its rate of change? then why does it say displacement? that is the same thing right?

zepdrix (zepdrix):

We're taking the idea of rate of change, and giving it a direct application. It's important because displacement, velocity, acceleration have a very clear connection to one another so it's important to get comfortable with the notation. Especially if you ever plan on taking a physics course. :)

zepdrix (zepdrix):

Do you understand how to calculate s(3) and s(1)? :)

OpenStudy (laddiusmaximus):

not really

OpenStudy (laddiusmaximus):

its not just 3-1 is it?

zepdrix (zepdrix):

\[\large s(t)=1+2t+\frac{ 1 }{ 4 }t^2\] \[s(3)=1+2(3)+\frac{ 1 }{ 4 }(3)^2\] Understand? :) Try to find s(3) and s(1).

OpenStudy (laddiusmaximus):

hmmm I think I tried that once and thought it was wrong

zepdrix (zepdrix):

Hmmm :o

OpenStudy (laddiusmaximus):

well for s(3) I got 37/4

OpenStudy (laddiusmaximus):

s(1)=13/4

zepdrix (zepdrix):

k looks good. plug them bad boys in!

OpenStudy (laddiusmaximus):

wait what do I do with those afterwards? I do that for all the intervals?

zepdrix (zepdrix):

\[\large v_{ave}=\frac{ s(3)-s(1) }{ 3-1 }\] You throw them into this formula to find the average velocity between time 1 and 3 seconds.

OpenStudy (laddiusmaximus):

crap right.

OpenStudy (laddiusmaximus):

so 37/4-13/4/37/4-13/4 ?

zepdrix (zepdrix):

the bottom is just 3 and 1.

OpenStudy (laddiusmaximus):

so I plugged them in and the velocity appears to be decreasing

zepdrix (zepdrix):

\[\large \frac{ s(3)-s(1) }{ 3-1 }=\frac{ \frac{ 37 }{ 4 }-\frac{ 13 }{ 4 } }{ 3-1 }=\frac{ \frac{ 24 }{ 4 } }{ 2 }=\frac{ 24 }{ 8 }=3\] Hmm did you come up with something like this?

OpenStudy (laddiusmaximus):

yup then 1, 42/16,1010/400

zepdrix (zepdrix):

Oh you did the others as well? neato.

OpenStudy (laddiusmaximus):

so what do I do with them? where does instantaneous velocity fit it? when I take the limit as x approaches 1?

zepdrix (zepdrix):

\[\large s'(t)=v(t)\]v(1)=instantaneous rate of change of velocity at t=1.

OpenStudy (laddiusmaximus):

so I take the derivative?

zepdrix (zepdrix):

mhm

OpenStudy (laddiusmaximus):

you know Im very tired. I cant think right now. Im going to tackle the rest of this tomorrow. thanks for all your help.

zepdrix (zepdrix):

hehe ok np c: haf good nite